OpenStudy (anonymous):

Which rectangular equation is equivalent to the polar equation r csc theta = 8 ? x = 8 y = 8 x = 8y^2

OpenStudy (anonymous):

or x^2 + (y-4)^2 = 16

OpenStudy (anonymous):

@theEric

OpenStudy (anonymous):

I think its first or last option?

OpenStudy (theeric):

So, \(r\csc(\theta)\tan(\theta)=8\) ?

OpenStudy (theeric):

This is a bunch of algebra!

OpenStudy (anonymous):

yeah I'm starting to notice that

OpenStudy (theeric):

What would be a good first step? I haven't gotten to any of the given answers, but I do have a rectangular equation.

OpenStudy (anonymous):